To what volume should you dilute 135mL of an 8.15M CuCl2 solution so that 48.0mL of the diluted solution contains 4.22g CuCl2?

Express your answer in milliliters.
Answer #1

moles CuCl2 = 5.7 g / 135.34 g/mol= 0.0421
M = 0.0421/ 0.0600 L= 0.702

0.155 x 8.80= V x 0.702

V = 1.94 L => 1940 mL